### v1

# DiaHalu: A Dialogue-level Hallucination Evaluation Benchmark for Large Language Models

Abstract
Since large language models (LLMs) achieve significant success in recent years, the hallucination issue remains a challenge, numerous benchmarks are proposed to detect the hallucination. Nevertheless, some of these benchmarks are not naturally generated by LLMs but are intentionally induced. Also, many merely focus on the factuality hallucination while ignoring the faithfulness hallucination. Additionally, although dialogue pattern is more widely utilized in the era of LLMs, current benchmarks only concentrate on sentence-level and passage-level hallucination. In this study, we propose DiaHalu, the first dialogue-level hallucination evaluation benchmark to our knowledge. Initially, we integrate the collected topics into system prompts and facilitate a dialogue between two ChatGPT3.5. Subsequently, we manually modify the contents that do not adhere to human language conventions and then have LLMs re-generate, simulating authentic human-machine interaction scenarios. Finally, professional scholars annotate all the samples in the dataset. DiaHalu covers four common multi-turn dialogue domains and five hallucination subtypes, extended from factuality and faithfulness hallucination. Experiments through some well-known LLMs and detection methods on the dataset show that DiaHalu is a challenging benchmark, holding significant value for further research.
DiaHalu: A Dialogue-level Hallucination Evaluation Benchmark for Large Language Models
Kedi Chen, Qin Chenthanks: Corresponding author., Jie Zhou, Yishen He and Liang He School of Computer Science and Technology, East China Normal University, Shanghai, China {kdchen,10215102502}@stu.ecnu.edu.cn {qchen, jzhou, lhe}@cs.ecnu.edu.cn
1 Introduction
Figure 1: Our benchmark not only includes factuality hallucination but also incorporates faithfulness hallucination at the dialogue level, although most benchmarks overlook the latter.
Large language models (LLMs) (Zhao et al., 2023a), addressing many downstream tasks through natural language generation (NLG) technique, achieve significant success across diverse fields of natural language processing (NLP) (Li et al., 2024; Pan et al., 2024; Wei et al., 2023; Fan et al., 2023). With a considerable volume of parameters and sophisticated training methodologies, LLMs significantly propelled advancements in artificial intelligence (Zhao et al., 2023a).
Despite many advantages of large language models, the issue of hallucination remains a primary concern (Ji et al., 2023; Zhang et al., 2023). Hallucination predominantly centers on the inclination of LLMs to generate nonsensical or untruthful contents for specific sources (Wang et al., 2023). Therefore, the occurrence of hallucinations poses risks to the application of large language models in various real-world scenarios (Omiye et al., 2023; Kasneci et al., 2023; Wu et al., 2023a).
Given the aforementioned risks, hallucination detection emerges as a highly crucial task. In recent years, researchers propose numerous benchmarks for hallucination detection task (Li et al., 2023; Guan et al., 2023; Manakul et al., 2023; Yang et al., 2023). Nevertheless, several problems persist in these benchmarks. (1) Not naturally generated. One pitfall of existing benchmarks for detecting LLMs’ hallucination is that the hallucinated contents are typically induced via manually designed trigger prompts (Li et al., 2023), while not naturally generated by LLMs as in daily usage (Liu et al., 2022). (2) Merely focusing on factuality hallucination. Most benchmarks merely focus on detecting factuality hallucination (Guan et al., 2023), with few datasets that can demonstrate faithfulness hallucination (Huang et al., 2023a) (Figure 1). (3) Only concentrating on sentence-level and passage-level. Researchers proposes many sentence-level (Manakul et al., 2023; Zhao et al., 2023b) and passage-level (Yang et al., 2023; Feng et al., 2023b) hallucination detection benchmarks. However, the dialogue pattern has broader and more widespread applications in LLMs. Although it is not mentioned in previous researches to our current knowledge, dialogue-level hallucination detection is equally essential.
Therefore, we propose a new dialogue-level hallucination evaluation benchmark for large language models (DiaHalu). We initially determine four domains for multi-turn dialogue: knowledge-grounded, task-oriented, chit-chat and reasoning. For each domain, we undertake a three-step process to construct the dataset. (1) We collect topics for dialogues from various sources. (2) We incorporate the topics into artificially designed system prompts and input them into two ChatGPT3.5, enabling them to engage in a multi-turn dialogue. (3) Since the knowledge-grounded and task-oriented domains stand for real human-machine interaction scenarios, we align the contents of one of the conversational participants with human language. We manually modify the contents that do not conform to human language conventions and have LLMs re-generate, resulting in the final outcomes. Then, professional scholars annotate all the samples with labels, hallucination types and locations, as well as explanations. It is noteworthy that we not only consider the factuality hallucination but also further classify the faithfulness hallucination into three types: Incoherence, Irrelevance and Overreliance. We similarly introduce the reasoning hallucination for the reasoning domain. Additionally, we conduct experiments on DiaHalu by deploying existing hallucination detection methods and some commonly used LLMs. The results indicate that DiaHalu is a highly challenging dataset, and both Chain-of-Thought and retrieval methods can improve the performance of the detection. Our contributions can be listed as follows:
· To our current knowledge, we propose the first dialogue-level hallucination detection benchmark named DiaHalu [^1] https://anonymous.4open.science/r/DiaHalu-A1E9/.
· DiaHalu covers four multi-turn dialogue domains along with five hallucination subtypes extended from factuality and faithfulness hallucinations, which are more widely applicable in real-world scenarios.
· The experimental results indicate that DiaHalu is a highly challenging task for most LLMs and existing detection methods, holding significant value for further research.
2 Related Work
2.1 Hallucination Detection Benchmarks
Hallucination predominantly centers on the inclination of large language models to generate nonsensical or untruthful contents for specific sources (Wang et al., 2023). Therefore, detecting hallucination in LLMs is of great importance. In recent years, researchers propose numerous benchmarks for the hallucination detection task.
In earlier years, hallucination detection benchmarks are primarily curated through manual methods or generated via conventional language models. FactCollect (Ribeiro et al., 2022) is an artificially generated, multi-source factual hallucination detection dataset. Muhlgay et al. (2023) collects error samples by instructing the language model based on pre-defined error types. HADES (Liu et al., 2022) constitutes a hallucination detection dataset produced by conventional language model BERT (Devlin et al., 2019). These benchmarks are not naturally generated by LLMs as in daily usage.
Consequently, some benchmarks are proposed to investigate the capacity for the direct generation of large language models. Zhao et al. (2023b); Fu et al. (2023); Chen et al. (2023a); Huang et al. (2023b); Zheng et al. (2023) enable LLMs to handle Question-Answer (QA) tasks and assess the factual accuracy of their responses. Concept-7 dataset used by Luo et al. (2023), evaluates whether a language model genuinely comprehends the meaning of each concept, thereby determining the presence of hallucination. FactCHD (Chen et al., 2023c) is generated based on natural language text and knowledge graphs (KGs), deploying two major steps: Query and Response Data Simulation & Chain of Evidence and Human Filter, to detect fact-conflicting hallucination in LLMs. Mündler et al. (2023) employs a generative language model (gLM) to rewrite sentences according to the given context. New sentences compose a dataset that can evaluate whether the generated sentences exhibit knowledge-based self-contradiction hallucination. The aforementioned benchmarks mainly focus on detecting factuality hallucination (Guan et al., 2023), while ignoring the faithfulness hallucination (Huang et al., 2023a). The benchmark proposed in this paper extends to include faithfulness hallucination, that is, to evaluate the coherence and relevance of generated contents by LLMs.
Researchers also raise many sentence-level (Manakul et al., 2023; Zhao et al., 2023b; Guan et al., 2023) and passage-level (Yang et al., 2023; Feng et al., 2023b; Li et al., 2023) hallucination detection benchmarks. Nevertheless, the dialogue pattern holds broader applications within LLMs. While previous research does not, to our current knowledge, propose a dialogue-level hallucination detection benchmark. So, our DiaHalu is at the dialogue level.
2.2 Dialogue System in LLMs
It is the formidable dialogue capability exhibited by ChatGPT3.5 (Wu et al., 2023b) and GPT4 (OpenAI, 2023), that stimulates a heightened interest among researchers in large language models. The ability for multi-turn dialogue is among the most crucial features of LLMs, which has broader and more widespread applications in the current development of NLP.
Dialogue system can be primarily categorized into various types, including task-oriented dialogue (Wen et al., 2017), chit-chat dialogue (Sun et al., 2021), knowledge-grounded dialogue (Ghazvininejad et al., 2018), etc. In these scenarios, language models are required to exhibit long-term memory, recognize shifts in dialogue states, encompass extensive knowledge, and engage in reasoning processes (Huang and Chang, 2023). Despite the remarkable general capability of LLMs, there are still deficiencies in handling complex dialogue scenarios (Chen et al., 2017; Deriu et al., 2021).
Due to the critical significance of the dialogue ability, this paper underscores the detection of hallucination produced by LLMs in multi-turn dialogues.
Figure 2: The demonstration of the DiaHalu benchmark, which covers four domains and five hallucination types within dialogue-level scenarios. We also provide explanations and sources in the benchmark.
3 The Overview of DiaHalu
3.1 Principles
The primary objective of this benchmark is to conduct hallucination detection in large language models. Hence, it is imperative to comprehend the meaning of hallucination. Hallucination predominantly centers on the inclination of LLMs to generate nonsensical or untruthful contents regarding specific sources (Wang et al., 2023). The significance of hallucination detection lies in elevating the quality of text generation, preventing misleading information and misunderstandings, supporting applications within professional domains, etc. Therefore, to enhance the universality of our benchmark, it encompasses various multi-turn dialogue scenarios and multiple types of hallucination, extending from factuality hallucination and faithfulness hallucination (Huang et al., 2023a).
3.2 Hallucination on Diverse Domains
We consider hallucination in diverse domains of multi-turn dialogue scenarios. Our benchmark covers a total of four domains, which are shown in Figure 2. Below are their specific descriptions.
Knowledge-grounded dialogue
is designed for users to engage in knowledge-based dialogue with LLMs (Ghazvininejad et al., 2018). The knowledge includes world knowledge, factual knowledge, commonsense knowledge, and multi-hop web knowledge. It principally examines the accuracy of knowledge of the parameters in LLMs (Petroni et al., 2019).
Task-oriented dialogue
is in a form of human-computer interaction, intending to accomplish a user-specified task (Wen et al., 2017). This type of dialogue system focuses on understanding the users’ task requirements and utilizes a LLM to provide relevant information or perform specific tasks accordingly.
Chit-Chat dialogue
involves open-ended and non-goal dialogues (Sun et al., 2021). We provide two LLMs with personas and facilitate a dialogue between them. This approach allows for the evaluation of their memory capabilities, conversational coherence, and relevance to the information being discussed.
Reasoning dialogue
centralizes on the logical reasoning and understanding capabilities of LLMs. We have the models discuss mathematical problems to achieve the answers (Kakarla et al., 2024).
3.3 Hallucination Types
We consider factuality hallucination and faithfulness hallucination (Huang et al., 2023a). Based on Chen et al. (2023b), we carry out a detailed classification of faithfulness hallucination into Incoherence, Irrelevance, and Overreliance. Meanwhile, we introduce Reasoning Error within the reasoning dialogue. The example of each type can be referenced in Figure 2 and Appendix A.4.
Non-factual
implies whether it aligns with factual information.
Incoherence
includes input-conflicting, context-conflicting and self-conflicting contents in the dialogue.
Irrelevance
means that something unrelated to the topic of the conversation comes up.
Overreliance
is that the LLM excessively relies on the correctness of context, generating responses for unanswerable content (Slobodkin et al., 2023).
Reasoning Error
covers all errors within the reasoning dialogue.
4 The Construction of DiaHalu
4.1 The Collection of Dialogue Topics
Since we confirm four domains for DiaHalu, the first step is to collect the topics for each dialogic domain.
For knowledge-grounded dialogue, we take into account world knowledge, factual knowledge, commonsense knowledge and multi-hop web knowledge. Therefore, we gather dialogue topics for TruthfulQA (Lin et al., 2022), CommonsenseQA (Talmor et al., 2019) and CWQ (Talmor and Berant, 2018) datasets. There are also topics provided by GPT4 (OpenAI, 2023) and social media (including the authors). As for task-oriented dialogue, we primarily apply the most widely used MultiWOZ (MultiWOZ 2.1) (Budzianowski et al., 2018) which has 7 real-life scenarios. To enrich the dialogue settings, we also consider the DSTC (DSTC 1.0) (Williams et al., 2013) dataset with a focus on bus routes. GPT4 and social media are harnessed to augment user behaviors and generate more dialogue occasions. We define the LLMs with personas primarily from Jandaghi et al. (2023) and facilitate an open chit-chat dialogue between them. Additionally, we make use of mathematical problems to assess the logical reasoning dialogue ability of LLMs. These problems are sourced from GSM8K (Cobbe et al., 2021) and MathQA (Amini et al., 2019), both of which involve mathematical problems and solving processes encountered by middle school students. The overall distribution of the above sources for dialogue topics is illustrated in Figure 3.
Figure 3: The distribution of the 10 sources for dialogue topics.
4.2 Dialogue Generation
Once finishing collecting the dialogue topics for each domain, we leverage ChatGPT3.5 to generate conversations in the format of self-dialogue. The complete process of dialogue generation is illustrated in Figure 4.
Figure 4: The complete process of dialogue generation.
Initially, we integrate the dialogue topics into two system prompts, which are then inputted separately into two ChatGPT3.5. These two system prompts guide the LLMs to generate N rounds of dialogue in a given domain and topic. More details of the system prompts can be found in Appendix A.1. Then, for knowledge-grounded dialogue and task-oriented dialogue, we manually examine all responses from A to determine their adherence to human language. For in both scenarios, we consider real human-machine interaction, aiming to assess the LLMs’ adaptability to genuine human behaviors. (We assume that A is the user and B is the LLM. In this setup, we ensure the accuracy of A and only annotate the contents of B.) The chit-chat dialogue and reasoning dialogue are relatively unconstrained and freely conducted, without incorporating any specific human-machine interaction settings (Section 3.2). They necessitate only their memory and comprehension capabilities regarding contextual information, thereby minimizing the need for manual intervention. Consequently, for the responses of A in knowledge-grounded and task-oriented scenarios, where the contents do not conform to human language, we manually modify and have LLMs re-generate. Eventually, we obtain the complete dataset of multi-turn dialogues.
We provide one output sample in Appendix A.2. The comparison with other benchmarks is in Appendix A.3, highlighting the distinctive features that our DiaHalu is a naturally generated dialogue-level benchmark by LLMs, with various kinds of hallucination and explanations. The statistical information of the whole benchmark is in Table 1.
Attribute Attribute Value
Benchmark Name DiaHalu
Generated by ChatGPT3.5
Sample Form dialogue-level
Sample Numbers 748
Dialogue Rounds 6-10
Avg. Rounds 6.8951
Domain Numbers 4
Hallucination Types 5
Max. Response Length (Words) 183
Avg. Response Length (Words) 13.5091
Table 1: The statistical information of the benchmark.
4.3 Human Annotation
Annotating the hallucination and its types in the benchmark is a very challenging task. Since there may be more than one instance of hallucination in multi-turn dialogues. Also, some hallucination types in edge cases are difficult to differentiate. Therefore, the entire annotation process demands a high level of expertise from annotators and requires detailed definitions for ambiguous contents.
The annotators
of our benchmark are all seasoned researchers in the field of linguistics and natural language processing. The annotators all obtain at least a bachelor’s degree, pass the IELTS or TOEFL exams, and are proficient in using search engines such as Google and Bing. In addition, everyone exhibits strong collaborative and communicative skills. To ensure the validity of the annotation information, we invite experienced experts in the field of LLMs’ hallucination detection from both academia and industry to conduct random sampling checks.
Annotation process
is divided into three steps. (1) Each annotator labels 50 samples for each domain, followed by a careful discussion. The annotators are required to label the presence of hallucination, hallucination types and locations, along with the corresponding explanations. The discussions intricately define the application scope of each hallucination label (Appendix A.4), and ultimately we decide to annotate the corresponding hallucination type on the first dialogue round in which hallucination occurs. (2) All the annotators label the entire dataset, discussions and corrections are made for inconsistent annotations. (3) Statistical analysis is performed on the annotated results. For more details, please refer to Appendix A.5.
Annotation Consistency
For evaluating the inter-annotator consistency, we calculate the Fleiss’s Kappa (Randolph, 2005) of Inter-Annotator Agreement (IAA) (Artstein, 2017), which is a statistical measure used to assess the degree of agreement among multiple raters for a set of items.
After the whole annotation process, we achieve a label matrix $\mathbf{L}\in\mathbb{R}^{N_{s}*N_{A}}$. $N_{s},N_{A}$ represent the number of dialogue samples and annotators respectively. The calculation for Fleiss’s Kappa is shown below:
$\begin{split}P_{e}&=\left(\frac{\sum_{i=1}^{N_{s}}\sum_{j=1}^{N_{A}}\chi_{\left\{0\right\}}(\mathbf{L}\left[i,j\right])}{N_{s}*N_{A}}\right)^{2}\\ &+\left(\frac{\sum_{i=1}^{N_{s}}\sum_{j=1}^{N_{A}}\chi_{\left\{1\right\}}(\mathbf{L}\left[i,j\right])}{N_{s}*N_{A}}\right)^{2}\end{split}$ (1)
$\begin{split}P_{o}&=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\frac{\left(\sum_{j}^{N_{A}}\chi_{\left\{0\right\}}(\mathbf{L}\left[i,j\right])\right)^{2}}{{N_{A}*\left(N_{A}-1\right)}}\\ &+\frac{\left(\sum_{j}^{N_{A}}\chi_{\left\{1\right\}}(\mathbf{L}\left[i,j\right])\right)^{2}-N_{A}}{N_{A}*\left(N_{A}-1\right)}\end{split}$ (2)
$Fleiss^{\prime}sKappa=\frac{P_{o}-P_{e}}{1-P_{e}}$ (3)
where $P_{o}$ and $P_{e}$ represent the relative observed agreement among annotators and the hypothetical probability of chance agreement respectively. $\chi_{y}(Y)$ is the Indicator Function, which means when the value of $Y$ is in set $y$, the whole function equals 1.
According to the above formulas, the calculated result for Fleiss’s Kappa of our benchmark is 0.8709, representing almost perfect agreement among all the annotators.
Annotation Results
After annotating the entire dataset, we conduct several statistical analyses on it. Table 2 reveals the probability of hallucination occurring in each dialogue domain. The results indicate that hallucinations are highly likely to arise in Knowledge-grounded dialogue and reasoning dialogue. Therefore, the knowledge and reasoning abilities of LLMs still need further improvement. Despite LLMs’ powerful multi-turn dialogue capability, faithfulness hallucination such as irrelevance, incoherence and overreliance still persists. Figure 5 presents the proportion of each hallucination type in each dialogue domain. Irrelevance, incoherence, and overreliance Widely exit in daily dialogue contexts, such as task-oriented and chit-chat scenarios. In knowledge-grounded dialogue, the factuality hallucination constitutes a significant proportion, while in reasoning dialogue, almost all hallucinations are defined as errors in reasoning. This statistical information can help us understand the types of hallucination in LLMs’s multi-turn dialogues, facilitating an exploration of their origins and contributing to the elimination of these types of hallucination.
Knowledge Task Chit Reasoning Overall
# Number
272
131
186
159
748
# Non-Halu
141
84
115
79
419
# Halu
131
47
71
80
329
Halu Rate (%)
48.16
35.88
38.17
50.31
43.98
Table 2: The statistical information of hallucination on the four dialogue domains.
Figure 5: The distribution of five different hallucination types within the four dialogue domains.
Method Knowledge-grounded Task-oriented Chit-Chat Reasoning Overall
Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1
Random 48.19 30.53 37.38 35.71 21.28 26.67 32.50 18.31 23.42 50.00 23.75 32.20 43.39 24.92 31.66
$SelfCheckGPT_{B}$ 47.92 17.56 25.70 32.14 19.15 24.00 37.93 15.49 22.00 57.50 28.75 38.33 45.52 20.06 27.85
$SelfCheckGPT_{N}$ 64.41 29.01 40.00 34.25 53.19 41.67 46.58 47.89 47.22 64.00 20.00 30.48 49.13 34.35 40.43
$SelfCheckGPT_{P}$ 56.45 26.72 36.27 44.73 36.17 40.00 43.33 54.92 48.45 60.96 62.50 61.73 51,84 42.86 46.92
LLaMA-30B 38.46 3.82 6.94 40.00 4.26 7.69 50.00 11.27 18.39 71.43 6.25 11.49 48.78 6.08 10.81
Vicuna-33B 50.00 4.58 8.39 100.00 4.26 8.16 28.57 2.82 5.13 41.38 15.00 22.02 44.00 6.69 11.61
ChatGPT3.5 50.00 0.76 1.50 0.00 0.00 0.00 42.86 4.23 7.69 55.56 6.25 11.24 45.00 2.74 5.16
ChatGPT4 82.46 35.88 50.00 73.68 29.79 42.42 56.52 18.31 27.66 71.43 81.25 76.02 73.16 42.25 53.56
Table 3: The classification results on four kinds of baselines, and the best F1 scores are in bold form.
Method Knowledge-grounded Task-oriented Chit-Chat Reasoning Overall
Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1
ChatGPT3.5 50.00 0.76 1.50 0.00 0.00 0.00 42.86 4.23 7.69 55.56 6.25 11.24 45.00 2.74 5.16
w/ CoT 45.45 3.82 7.04 20.00 2.13 3.85 28.57 2.82 5.13 53.85 8.75 15.05 41.67 4.56 8.22
w/ retrieval 70.00 5.34 9.93 - - - - - - 71.43 12.50 21.27 - - -
ChatGPT4 82.46 35.88 50.00 73.68 29.79 42.42 56.52 18.31 27.66 71.43 81.25 76.02 73.16 42.25 53.56
w/ CoT 82.02 55.73 66.36 72.41 44.68 55.26 69.05 40.85 51.33 69.52 91.25 78.92 73.96 59.57 65.99
w/ retrieval 75.90 48.09 58.88 - - - - - - 70.65 81.25 75.58 - - -
Table 4: The results of CoT and retrieval techniques which can enhance hallucination detection on ChatGPT3.5 and ChatGPT4.
5 Experiments
In this section, we assess the performance of several evaluation models and methods on the dataset we introduced. Thereby, we can determine the effectiveness of existing methods in detecting dialogue-level hallucination.
5.1 Baselines
We select some powerful LLMs by providing specific prompts, to assess the presence of hallucinations. Similarly, we also conduct experiments using existing hallucination detection methods.
Random
A straightforward approach that randomly generates a label for each sample.
SelfCheckGPT
(Manakul et al., 2023) It is a widely used black-box hallucination detection framework. It rephrases the contents to be detected while ensuring the consistency of semantics by LLMs with different temperatures. Furthermore, it calculates the consistency between the original and the rephrased contents using five methods, thereby determining whether there exits hallucination. The indices $B$, $N$ and $P$ respectively denote scoring with BERTScore (Zhang et al., 2020), scoring with Natural Language Inference methods (He et al., 2023) and the direct judgment using prompts.
LLaMA-30B && Vicuna-33B
They are two well-pretrained and widely deployed open-source LLM backbones [^2] https://huggingface.co/huggyllama/llama-30b, https://huggingface.co/lmsys/vicuna-33b-v1.3 (Touvron et al., 2023; Chiang et al., 2023). We provide a specially designed prompt to assist with detection. More details are shown in Appendix A.6.
ChatGPT3.5 && ChatGPT4
Both of these models are developed by OpenAI [^3] https://chat.openai.com/. ChatGPT3.5 marks the beginning of the era of large language models and GPT4 is currently the most powerful language model (Wu et al., 2023b; OpenAI, 2023). We employ the same prompt as LLaMA and Vicuna do.
5.2 Metrics
Despite we annotating the types of hallucinations in the dataset, achieving consistent labels even among humans requires further discussion. Therefore, similar to past hallucination detection efforts, we merely focus on a binary classification task of determining the existence of hallucination. Consequently, we utilize binary classification evaluation metrics. The positive label for this classification task is set as "Halu", for our main focus is testing the model’s ability to recognize hallucination.
5.3 Main Results
From the results in Table 3, we can get the following conclusions.
First, DiaHalu is a very challenging benchmark for dialogue-level hallucination detection.
Except for ChatGPT4, the F1 scores of all other detection methods do not exceed 50.00. Existing language models, such as LLaMA-30B and Vicuna-33B, are not effective in accurately discerning most samples that involve hallucinations. Regarding the specialized detection method SelfCheckGPT, it achieves relatively better performance when applying the prompt and NLI methods. However, distinguishing between positive and negative samples proves challenging with BERTScore.
Second, when performing self-detection using ChatGPT3.5, there is a noticeable phenomenon of overconfidence.
Our dataset is primarily generated by ChatGPT3.5, which exhibits high confidence in its output. Despite providing a specially designed detection prompt, it still struggles to differentiate whether the dialogue content is hallucinating or not. So, the majority of its output labels are "Non-Halu".
Third, the faithfulness hallucination is more difficult to detect for LLMs.
Apart from the specialized hallucination detection method SelfCheckGPT, the results from directly harnessing LLMs for judgment indicate that the recognition accuracy for task-oriented and chit-chat domains of dialogue are much lower than that for the knowledge-grounded and reasoning dialogue. This is because the hallucination types in the knowledge-grounded and reasoning dialogue are primarily Non-factual and Reasoning Error, which present in a more direct and apparent manner. While task-oriented and chit-chat domains mainly consist of three types of faithfulness hallucination, which requires a LLM to possess long-term memory and the ability to recognize topic transitions in dialogues.
5.4 Chain-of-Thought and Retrieval for Detection
Chain-of-Thought and Retrieval are two important techniques for enhancing the reasoning ability of LLMs. Chain-of-Thought (CoT) (Feng et al., 2023a) describes the organized sequence of logical reasoning that unfolds during thinking. Retrieval (Gao et al., 2023) means retrieving relevant contents from the media to supplement external knowledge for LLMs. In this section, we test whether these two techniques can improve the effectiveness of hallucination detection in Table 4. More details are depicted in Appendix A.7.
The experimental results indicate that both methods have facilitating effects on hallucination detection. However, the combination of GPT4 with retrieval methods shows a decrease of around 0.5 points in F1 score in the domain of reasoning detection. We believe that GPT4 inherently possesses a strong reasoning ability, and therefore, additional retrieval contents introduce noise to its judgments.
6 Conclusion
In this paper, we propose a dialogue-level hallucination evaluation benchmark. We construct the benchmark in a three-step process and have the experts annotate all the data. DiaHalu covers four multi-turn dialogue domains and five hallucination subtypes. Experiments through some well-known LLMs and detection methods on the dataset show that DiaHalu is a challenging task, holding significant value for further research (Appendix A.8).
Limitations
This paper proposes a novel dialogue-level hallucination detection evaluation benchmark named DiaHalu. The benchmark covers four multi-turn dialogue domains and five hallucination subtypes. There is significant value for further research. However, two main limitations also exist. (1) During the second step of the benchmark construction phase, aligning the contents of speaker A with human language consumes a considerable amount of time and effort. Frequent calls to the ChatGPT3.5 API_Keys result in a significant expense. Simultaneously, achieving consistency among all annotators led to prolonged discussion time. (2) We do not partition the dataset into training, validation, and test sets. This is because we consider the black-box detection scenario for those closed-source LLMs. If a division into these three types of datasets is required, it would necessitate more data samples and larger resource consumption.
Ethics Statement
The benchmark is primarily generated by ChatGPT3.5. We obtain all the API_Keys through a paid subscription. All the annotators are real and they receive corresponding compensation and rewards. The entire process and outcomes are free from intellectual property and ethical legal disputes.
Acknowledgements
Due to the constrictions of the double-blind review mechanism, we will acknowledge all contributors and institutions as soon as our paper is accepted.
Appendix A Appendices
A.1 System Prompts for Dialogue Generation
In this section, we present the specific form of the system prompts for the four dialogue domains. We use ChatGPT3.5 (gpt-3.5-turbo-1106) [^4] https://platform.openai.com/docs/guides/text-generation/chat-completions-api with the temperature 0.1. The brief system prompts of the knowledge-grounded dialogue, task-oriented dialogue, chit-chat dialogue and reasoning dialogue are respectively presented in Figure 6, Figure 7, Figure 8 and Figure 9.
Figure 6: The brief system prompts for knowledge-grounded dialogue.
Figure 7: The brief system prompts for task-oriented dialogue.
Figure 8: The brief system prompts for chit-chat dialogue.
Figure 9: The brief system prompts for reasoning dialogue.
A.2 Data Format
We provide the specific format of one sample from the benchmark in Figure 10.
Figure 10: One sample from the benchmark
A.3 The Comparison with Other Benchmarks
In Table 5, we present the differences between DiaHalu and other hallucination detection benchmarks, highlighting the distinctive features that our DiaHalu is a naturally generated dialogue-level benchmark by LLMs, with various kinds of hallucination and explanations. All the compared benchmarks can be referred to in Section 2.1. HaluEval, WikiBio+, and PHD benchmark come from the paper Li et al. (2023), Manakul et al. (2023) and Yang et al. (2023) respectively.
A.4 The Application Scope of Hallucination Labels
Annotating the hallucination and its types in the benchmark is a very challenging task. One of the reasons is that some hallucination types in side cases are defiant to differentiate. Therefore, in the first stage of annotation, we provide detailed definitions for each hallucination type.
Non-factual
implies that it does not align with facts or introduce elements that do not exist in real life.
Incoherence
means there is a contradiction between the dialogue role and his context, a contradiction with the dialogue object, or an internal contradiction within the current response (in a declarative sentence). It also refers to factual and relevant nonsense or contextually inappropriate responses.
Irrelevance
involves responses that are irrelevant to the dialogue topic or are unrelated due to a misunderstanding (in the case of a question).
Overreliance
is that the LLM excessively relies on the correctness of context, generating responses for unanswerable content (in a declarative sentence).
Reasoning Error
covers all errors within the reasoning dialogue.
Benchmark By LLMs Faith Halu Dialogue Explanation
FactCollect - - - -
HADES - $\checkmark$ - $\checkmark$
FactCHD $\checkmark$ - - $\checkmark$
HaluEval - - $\checkmark$ $\checkmark$
WikiBio+ $\checkmark$ - - -
PHD $\checkmark$ - - -
Ours $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 5: The comparison between our DiaHalu and other benchmarks. ‘By LLMs’, ‘Faith Halu’, ‘Dialogue’, and ‘Explanation’ mean whether it is naturally generated by LLMs, whether it provides faithfulness hallucination, whether it is at dialogue level, and whether there are explanations respectively.
A.5 The Supplementary Details for Annotation
We use Label Studio  [^5] https://labelstud.io/ for labeling, which is an online open-source data labeling platform in the field of artificial intelligence. The annotation interface is depicted in Figure 11.
Figure 11: The annotation interface on Label Studio.
A.6 The Prompt Designed for Detection
In Figure 12, we show the whole prompt specially designed for hallucination detection of the baselines: LLaMA-30B, Vicuna-33B, ChatGPT3.5, and ChatGPT4. It is worth noting that due to the poor instruction-following ability and the disorderly output format of the LLaMA and Vicuna models, we conduct experiments in a 1-shot manner. The ChatGPT3.5 used for detection is ChatGPT3.5 (gpt-3.5-turbo-1106) [^6] https://platform.openai.com/docs/guides/text-generation/chat-completions-api and the ChatGPT4 is ChatGPT4 (gpt-4-1106-preview)  [^7] https://openai.com/gpt-4.
Figure 12: The whole prompt for hallucination detection of the baselines.
A.7 The Settings for CoT and Retrieval
We employ CoT in all four domains to enhance the performance of ChatGPT3.5 and ChatGPT4. The specific CoT is illustrated in Figure 13. While only knowledge-grounded and reasoning domains are tested with retrieval via Google [^8] https://console.cloud.google.com/apis/library. This is because domains of task-oriented and chit-chat mainly involve scenarios related to daily life or virtual worlds, without specific domain knowledge as supplementary information.
Figure 13: The whole CoT for the four domains of dialogue.
A.8 Future Works
Dialogue-level hallucination detection
is an important work in the future. We propose the first dialogue-level hallucination detection evaluation benchmark and experimental results show that it is a very challenging task. Therefore, developing methods based on this dataset to achieve a relatively high recognition accuracy is highly valuable.
Dialogue-level hallucination elimination
is an extension task of this work. Most existing hallucination elimination methods primarily focus on passage-level factuality hallucination (Luo et al., 2024). Hallucination elimination at the dialogue level not only requires the models to have a long-context memory capability but also demands the abilities to recognize changes in topics and logical transitions in the dialogue. These abilities are also helpful in addressing faithfulness hallucinations. At the same time, improving the accuracy of knowledge in the LLM parameters and the reasoning ability are equally important.
Hallucination snowballing
is the phenomenon that LLMs tend to accumulate hallucinations rather than self-correcting during the generation process (Zhang et al., 2023). Some previous works also validate this phenomenon (Azaria and Mitchell, 2023; Ang et al., 2023). In our dataset, there is a noticeable issue of hallucination snowballing, so we only label the first occurrence of hallucination in our dataset. Eliminating hallucination snowballing is also important in the future.
The Unanswerability of LLMs
During the annotation process of this dataset, we introduce a hallucination category termed "overreliance", which represents answering unanswerable content (Slobodkin et al., 2023; Sulem et al., 2021). This phenomenon signifies that LLMs tend to trust the input provided by users. Sometimes, even when there are errors in user input, the LLMs still fail to recognize them. A few past researches explore the related areas. However, the issue in the application of human-machine interaction and multi-agent scenarios remains crucial.