### v2

# DiaHalu: A Dialogue-level Hallucination Evaluation Benchmark for Large Language Models

Abstract
Since large language models (LLMs) achieve significant success in recent years, the hallucination issue remains a challenge, and numerous benchmarks are proposed for hallucination detection. Nevertheless, some of these benchmarks are not naturally generated by LLMs but are intentionally induced. Also, many merely focus on the factuality hallucination while ignoring the faithfulness hallucination. Additionally, although dialogue pattern is more widely utilized in the era of LLMs, current benchmarks only concentrate on sentence-level and passage-level hallucination. In this study, we propose DiaHalu, the first dedicated dialogue-level hallucination evaluation benchmark for LLMs to our knowledge. Initially, we integrate the collected topics into system prompts and facilitate a dialogue between two LLMs. Subsequently, we manually modify the contents that do not adhere to human language conventions and then have LLMs re-generate, simulating authentic human-machine interaction scenarios. Finally, professional scholars annotate all the samples in the dataset. DiaHalu covers four common multi-turn dialogue domains and five hallucination subtypes, extended from factuality and faithfulness hallucination. Experiments through some well-known LLMs and detection methods on the dataset show that DiaHalu is a challenging benchmark, holding significant value for further research [^1] https://github.com/141forever/DiaHalu.
DiaHalu: A Dialogue-level Hallucination Evaluation Benchmark
for Large Language Models
Kedi Chen, Qin Chenthanks: Corresponding author., Jie Zhou, Yishen He  and Liang He School of Computer Science and Technology, East China Normal University, Shanghai, China {kdchen,10215102502}@stu.ecnu.edu.cn {qchen, jzhou, lhe}@cs.ecnu.edu.cn
1 Introduction
Figure 1: Our benchmark not only includes factuality hallucination but also incorporates faithfulness hallucination at the dialogue level, although most benchmarks overlook the latter one.
Large language models (LLMs) (Zhao et al., 2023b), addressing many downstream tasks through natural language generation (NLG) technique, achieve significant success across diverse fields of natural language processing (NLP) (Li et al., 2024a; Pan et al., 2024; Wei et al., 2023; Fan et al., 2023). With a considerable volume of parameters and sophisticated training methodologies, LLMs significantly propelled advancements in artificial intelligence (Zhao et al., 2023b).
Despite many advantages of large language models, the issue of hallucination remains a primary concern (Ji et al., 2023; Zhang et al., 2023c). Hallucination predominantly centers on the inclination of LLMs to generate nonsensical or untruthful contents for specific sources (Wang et al., 2023a). Therefore, the occurrence of hallucination poses risks to the application of large language models in various real-world scenarios (Omiye et al., 2023; Wu et al., 2023b; Chen et al., 2024).
Given the aforementioned risks, hallucination detection emerges as a highly crucial task. In recent years, researchers propose numerous benchmarks for hallucination detection task (Li et al., 2023b; Guan et al., 2023; Manakul et al., 2023; Yang et al., 2023). Nevertheless, several problems persist in these benchmarks. (1) Not naturally generated. One pitfall of existing benchmarks for detecting LLMs’ hallucination is that the hallucinated contents are typically induced via manually designed trigger prompts (Li et al., 2023b), while not naturally generated by LLMs as in daily usage (Liu et al., 2022). (2) Merely focusing on factuality hallucination. Most benchmarks merely focus on detecting factuality hallucination (Guan et al., 2023), with few datasets that can demonstrate faithfulness hallucination (Huang et al., 2023a) (Figure 1). (3) Only concentrating on sentence-level and passage-level. Researchers propose many sentence-level (Manakul et al., 2023; Zhao et al., 2023c) and passage-level (Yang et al., 2023; Feng et al., 2023b) hallucination detection benchmarks. However, the dialogue pattern has broader and more widespread applications in LLMs. Although it is rarely mentioned in previous researches, dialogue-level hallucination detection is equally essential.
Therefore, we propose a new dialogue-level hallucination evaluation benchmark for large language models (DiaHalu). We initially determine four domains for multi-turn dialogue: knowledge-grounded, task-oriented, chit-chat and reasoning. For each domain, we undertake a three-step process to construct the dataset. (1) We collect topics for dialogue from various sources, incorporate the topics into artificially designed system prompts and input them into two LLMs, enabling them to engage in a multi-turn dialogue. (2) Since the knowledge-grounded and task-oriented domains stand for real human-machine interaction scenarios, we align the contents of one of the conversational participants with human language. We manually modify the contents that do not conform to human language conventions and have LLMs re-generate, resulting in the final responses. (3) Professional scholars annotate all the samples with labels, hallucination subtypes and locations, as well as explanations. It is noteworthy that we not only consider the factuality hallucination but also further classify the faithfulness hallucination into three types: Incoherence, Irrelevance and Overreliance. We similarly introduce the reasoning hallucination for the reasoning domain. The advantages of ours compared with previous benchmarks are listed in Table 1. Additionally, we conduct experiments on DiaHalu by deploying existing hallucination detection methods and some commonly used LLMs. The results indicate that DiaHalu is a highly challenging benchmark. Our contributions can be listed as follows:
  • To our current knowledge, we propose the first dedicated dialogue-level hallucination detection benchmark for LLMs named DiaHalu.
  • DiaHalu covers four multi-turn dialogue domains along with five hallucination subtypes extended from factuality and faithfulness hallucination, which are more widely applicable in real-world scenarios.
  • The experimental results indicate that DiaHalu is a highly challenging benchmark for most LLMs and existing detection methods, holding significant value for further researches.
Benchmark By LLMs Faith Halu Multi Dia Explanation
FactCollect - - - -
BEGIN - - $\checkmark$ $\checkmark$
HADES - $\checkmark$ - $\checkmark$
FactCHD $\checkmark$ - - $\checkmark$
HaluEval - - $\checkmark$ $\checkmark$
WikiBio+ $\checkmark$ - - -
PHD $\checkmark$ - - -
Ours $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 1: The comparison between our DiaHalu and other benchmarks. ‘By LLMs’, ‘Faith Halu’, ‘Multi Dia’, and ‘Explanation’ mean whether it is naturally generated by LLMs, whether it provides faithfulness hallucination, whether it is at multi-turn dialogue level, and whether there are explanations respectively (Appendix A.1).
2 Related Work
2.1 Hallucination Detection Benchmarks
In recent years, researchers propose numerous benchmarks for hallucination detection.
In earlier years, hallucination detection benchmarks are primarily organized through manual methods or generated via conventional language models. FactCollect (Ribeiro et al., 2022) is an artificially generated, multi-source factual hallucination detection benchmark. Muhlgay et al. (2023) collects error samples by instructing the language model based on pre-defined error types. HADES (Liu et al., 2022) and BEGIN (Dziri et al., 2022c) constitute hallucination detection datasets by conventional language model BERT (Devlin et al., 2019) and T5 (mostly) (Raffel et al., 2020) respectively. These benchmarks are not naturally generated by LLMs as in daily usage.
Consequently, some benchmarks are proposed to investigate the direct generation abilities of large language models. Zhao et al. (2023c); Fu et al. (2023); Chen et al. (2023a); Huang et al. (2023b); Zheng et al. (2023) enable LLMs to handle Question-Answer (QA) task and assess the factual accuracy of their responses. Concept-7 dataset used by Luo et al. (2023) evaluates whether a language model truly comprehends the meaning of each concept, thereby determining the presence of hallucination. FactCHD (Chen et al., 2023c) is generated based on natural language text and knowledge graphs (KGs). Mündler et al. (2023) employs a generative language model (gLM) to rewrite sentences according to the given context. New sentences compose a dataset that can evaluate whether the generated sentences exhibit knowledge-based self-contradiction hallucination. The aforementioned benchmarks mainly focus on detecting factuality hallucination (Guan et al., 2023), while ignoring the faithfulness hallucination (Huang et al., 2023a). The benchmark proposed in this paper extends to include faithfulness hallucination, that is, to evaluate the coherence and relevance of contents generated by LLMs.
Researchers also raise many sentence-level (Manakul et al., 2023; Zhao et al., 2023c; Guan et al., 2023) and passage-level (Yang et al., 2023; Feng et al., 2023b; Li et al., 2023b) hallucination detection benchmarks. Nevertheless, the dialogue pattern holds broader applications within LLMs. While previous researches rarely, to our current knowledge, propose a dialogue-level hallucination detection benchmark for LLMs. So, our DiaHalu is at the dialogue level.
2.2 Hallucination Detection
Current methods for hallucination detection (Tonmoy et al., 2024) can mainly be divided into four categories. (1) Model-Based. This method involves having the language models perform a classification task to determine whether hallucinated contents are present (Zhao et al., 2023a; Maharaj et al., 2023). (2) Retrieval-Based. For the limited knowledge within the parameters of language models, we can detect hallucination by extracting or retrieving relevant knowledge from external knowledge graphs (Martino et al., 2023; Chen et al., 2023c) or web information sources (Béchard and Ayala, 2024). (3) Sample-Based. Another feasible method is to rewrite the generated contents to evaluate the consistency (Manakul et al., 2023; Zhao et al., 2023c; Zhang et al., 2023a). (4) Proxy-Based. The mainstream view of this method (Zhang et al., 2023b; Gupta et al., 2024) is that ‘the lower the probability of generating a token, the more likely a model is to produce hallucination’.
Figure 2: The demonstration of the DiaHalu benchmark, which covers four domains and five hallucination subtypes within dialogue-level scenarios. We also provide explanations and sources in the benchmark.
3 The Overview of DiaHalu
3.1 Principles
The primary objective of this benchmark is to conduct hallucination detection in large language models. Hence, it is imperative to comprehend the meaning of hallucination. Hallucination predominantly centers on the inclination of LLMs to generate nonsensical or untruthful contents regarding specific sources (Wang et al., 2023a). The significance of hallucination detection lies in elevating the quality of text generation, preventing misleading information and misunderstandings, supporting applications within professional domains, etc. Therefore, to enhance the universality of our benchmark, it encompasses various multi-turn dialogue scenarios and multiple subtypes of hallucination, extending from factuality hallucination and faithfulness hallucination (Huang et al., 2023a).
3.2 Hallucination on Diverse Domains
We consider hallucination in diverse domains of multi-turn dialogue scenarios. Our benchmark covers a total of four domains, shown in Figure 2. Their specific descriptions are in Appendix A.2.
Knowledge-grounded dialogue
is designed for users to engage in knowledge-based dialogue with LLMs (Ghazvininejad et al., 2018). The two speakers take part in a conversation about a knowledge-based issue.
Task-oriented dialogue
is in a form of human-computer interaction, intending to accomplish a user-specified task (Wen et al., 2017).
Chit-Chat dialogue
involves open-ended and non-goal dialogue (Sun et al., 2021). We provide two LLMs with personas and facilitate a dialogue between them.
Reasoning dialogue
Following previous works (Chen et al., 2023b; Li et al., 2023a; Buszydlik et al., 2024; Grover et al., 2024; Zheng et al., 2023; Huang and Chang, 2023), we also treat reasoning errors as a kind of hallucination. We have the models discuss mathematical problems to achieve the answers (Kakarla et al., 2024).
3.3 Hallucination Taxonomy
We consider both factuality and faithfulness hallucination (Huang et al., 2023a). Based on Chen et al. (2023b); Wu et al. (2023a); Dziri et al. (2022b) and early works on text coherence (Wolf and Gibson, 2004; Atwell et al., 2024), we carry out a detailed classification of faithfulness hallucination into Incoherence, Irrelevance, and Overreliance. Meanwhile, we introduce Reasoning Error within the reasoning dialogue (Chen et al., 2023b; Li et al., 2023a; Buszydlik et al., 2024; Grover et al., 2024; Zheng et al., 2023; Huang and Chang, 2023). The example of each type can be referenced in Figure 2.
Non-factual
implies whether it aligns with factual information.
Incoherence
includes input-conflicting, context-conflicting and self-conflicting contents in the dialogue.
Irrelevance
means that something unrelated to the topic of the conversation comes up.
Overreliance
is that the LLM excessively trusts in the correctness of the context, generating responses for unanswerable contents (Slobodkin et al., 2023).
Reasoning Error
covers all errors within the reasoning dialogue.
4 The Construction of DiaHalu
4.1 The Collection of Dialogue Topics
Since we confirm four domains for DiaHalu, the first step is to collect the topics for each dialogic domain.
For knowledge-grounded dialogue, we take into account world knowledge, factual knowledge, commonsense knowledge and multi-hop web knowledge. Therefore, we gather dialogue topics from TruthfulQA (Lin et al., 2022), CommonsenseQA (Talmor et al., 2019) and CWQ (Talmor and Berant, 2018) datasets. There are also topics provided by GPT4 (OpenAI, 2023) and social media (including the authors). As for task-oriented dialogue, we primarily apply the most widely used MultiWOZ (MultiWOZ 2.1) (Budzianowski et al., 2018) which covers 7 real-life scenarios. To enrich the dialogue settings, we also consider the DSTC (DSTC 1.0) (Williams et al., 2013) dataset with a focus on bus routes. GPT4 and social media are harnessed to augment user behaviors and generate more dialogue occasions. We define the LLMs with personas primarily from Jandaghi et al. (2023) and facilitate an open chit-chat dialogue between them. Additionally, we make use of mathematical problems to assess the logical reasoning dialogue abilities of LLMs. These problems are sourced from GSM8K (Cobbe et al., 2021) and MathQA (Amini et al., 2019), both of which involve mathematical problems and solving processes encountered by middle school students.
The overall distribution of the above sources for dialogue topics is illustrated in Appendix A.3.
4.2 Dialogue Generation
Once finishing collecting the dialogue topics for each domain, we leverage ChatGPT3.5 and GPT4 to generate conversations in the format of self-dialogue. The complete process of dialogue generation is illustrated in Figure 3.
Initially, we integrate the dialogue topics into two system prompts, which are then inputted separately into two LLMs (both are ChatGPT3.5 or GPT4). These two system prompts guide the LLMs to generate $N$ rounds of dialogue in a given domain and topic. More details of the system prompts can be found in Appendix A.4. Then, for knowledge-grounded dialogue and task-oriented dialogue, we manually examine all responses from A to determine their adherence to human language. For in both scenarios, we consider real human-machine interaction, aiming to assess the LLMs’ adaptability to genuine human behaviors. (We assume that A is the user and B is the LLM. In this setup, we ensure the accuracy of A and only annotate the contents of B.) The chit-chat dialogue and reasoning dialogue are relatively unconstrained and freely conducted, without incorporating any specific human-machine interaction settings (Section 3.2). They necessitate only their memory and comprehension capabilities regarding contextual information, thereby minimizing the need for manual intervention. Consequently, for the responses of A in knowledge-grounded and task-oriented scenarios, where the contents do not conform to human language, we manually modify and have LLMs re-generate. Eventually, we obtain the complete dataset of multi-turn dialogue.
Figure 3: The complete process of dialogue generation.
We provide one output sample and more generation details in Appendix A.5. The statistical information of the whole benchmark is in Table 2.
Attribute Attribute Value
Benchmark Name DiaHalu
Generated by ChatGPT3.5 / GPT4
Sample Form dialogue-level
Sample Numbers 1103
Dialogue Rounds 6-10
Avg. Rounds 6.9120
Domain Numbers 4
Hallucination Subtypes 5
Max. Response Length (Words) 183
Avg. Response Length (Words) 13.2899
Table 2: The statistical information of the benchmark.
4.3 Human Annotation
Annotating the hallucination and its types in this dataset is a very challenging task. Since there may be more than one instance of hallucination in multi-turn dialogue. Also, some hallucination subtypes in edge cases are difficult to differentiate. Therefore, the entire annotation process demands a high level of expertise from annotators and requires detailed definitions for ambiguous contents.
The annotators
of our dataset are all seasoned researchers in the field of linguistics and natural language processing. We invite experienced experts in the field of LLMs’ hallucination detection from both academia and industry to engage in discussion and conduct sampling checks. For more details about the annotators and the experts, refer to Appendix A.6 (The annotators).
Annotation process
is divided into three steps. (1) Each annotator labels some samples for each domain, followed by a careful discussion between the annotators and the experts. The discussions intricately define the application scope of each hallucination label (discussion results in Appendix A.7). (2) All the annotators label the entire dataset, discussions and corrections are made for inconsistent annotations. (3) Statistical analysis is performed on the annotated results. For more details, please refer to Appendix A.6 (The Annotation Process).
Annotation Consistency
For evaluating the inter-annotator consistency, we calculate the Fleiss’s Kappa (Randolph, 2005) of Inter-Annotator Agreement (IAA) (Artstein, 2017), which is a statistical measure used to assess the degree of agreement among multiple raters for a set of items. The final score of Fleiss’s Kappa is 0.8842, representing almost perfect agreement among all the annotators. For more calculation details, please refer to Appendix A.6 (Label Consistency).
Annotation Results
After annotating the entire dataset, we conduct several statistical analyses on it. Table 3 reveals the probability of hallucination occurring in each dialogue domain. The results indicate that hallucination are highly likely to arise in knowledge-grounded dialogue and reasoning dialogue. Therefore, the knowledge and reasoning abilities of LLMs still need further improvement. Despite LLMs’ powerful multi-turn dialogue capability, faithfulness hallucination such as irrelevance, incoherence and overreliance still persists. Figure 4 presents the proportion of each hallucination subtype in each dialogue domain. Irrelevance, incoherence, and overreliance widely exist in daily dialogue contexts, such as task-oriented and chit-chat scenarios. In knowledge-grounded dialogue, the factuality hallucination constitutes a significant proportion, while in reasoning dialogue, almost all hallucination are defined as errors in reasoning. This statistical information can help us understand the subtypes of hallucination in LLMs’ multi-turn dialogue, facilitating an exploration of their origins and contributing to the elimination of these subtypes of hallucination.
Knowledge Task Chit Reasoning Overall
# Number 371 210 263 259 1103
# Non-Halu 199 135 164 129 627
# Halu 172 75 99 130 476
Halu Rate (%) 46.36 35.71 37.64 50.19 43.16
Table 3: The statistical information of hallucination on the four dialogue domains. ‘# number’, ‘# Non-Halu’,‘# Halu’ and ‘Halu Rate’ represent the number of samples, the number of samples without hallucination, the number of samples with hallucination and the proportion of hallucinated samples.
Figure 4: The distribution of five different hallucination subtypes within the four dialogue domains.
Method Knowledge-grounded Task-oriented Chit-Chat Reasoning Overall
Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1
Random 41.57 43.02 42.29 31.86 48.00 38.30 38.46 50.51 43.67 49.61 49.23 49.42 40.72 47.06 43.66
$SelfCheckGPT_{B}$ 42.55 23.26 30.08 35.38 30.67 32.86 30.00 18.18 22.64 60.81 34.61 44.12 43.00 26.47 32.77
$SelfCheckGPT_{N}$ 59.46 25.58 35.77 38.84 62.67 47.96 45.19 47.47 46.30 70.58 18.46 29.27 48.65 34.03 40.05
$SelfCheckGPT_{P}$ 55.22 21.51 30.96 48.00 32.00 38.40 45.00 45.45 45.23 62.37 44.62 52.02 52.90 34.45 41.73
FOCUS 46.11 48.26 47.16 34.09 60.00 43.48 36.56 49.49 42.06 50.56 34.62 41.10 41.49 46.64 43.92
LLaMa-30B 37.50 5.23 9.18 30.77 5.33 9.09 50.00 11.11 18.18 81.25 10.00 17.81 49.33 7.78 13.43
Vicuna-33B 45.45 5.81 10.31 42.86 4.00 7.32 36.36 4.04 7.27 51.35 14.62 22.75 46.75 7.56 13.02
Gemini1.5 PRO 80.00 20.93 33.18 60.00 36.00 45.00 70.37 38.38 49.67 73.63 51.54 60.63 71.49 35.29 47.26
ChatGPT3.5 25.00 0.58 1.14 33.33 2.67 4.93 55.56 5.05 9.26 57.14 6.15 11.11 48.48 3.36 6.27
GPT4 80.89 31.98 45.83 74.19 30.67 43.40 67.74 21.21 32.31 74.07 61.54 67.23 75.21 37.61 50.14
Table 4: The classification results on four kinds of baselines, and the best F1 scores are in bold form. The indices $B$, $N$ and $P$ of SelfCheckGPT denote scoring with BERTScore, with NLI and using prompts, respectively.
Method Knowledge-grounded Task-oriented Chit-Chat Reasoning Overall
Prec Rec F1 Prec Rec F1 Prec Rec F1 Prec Rec F1 Prec Rec F1
FaithCritic w/retrieval 28.26 84.54 42.38 - - - - - - 51.63 79.17 62.50 - - -
Gemini1.5 PRO 80.00 20.93 33.18 60.00 36.00 45.00 70.37 38.38 49.67 73.63 51.54 60.63 71.49 35.29 47.26
w/ CoT 81.25 22.67 35.45$\uparrow$ 69.77 40.00 50.85$\uparrow$ 75.00 36.36 48.98$\downarrow$ 72.92 53.85 61.95$\uparrow$ 74.47 36.76 49.23$\uparrow$
w/ retrieval 86.04 21.51 34.42$\uparrow$ - - - - - - 76.70 60.77 67.81$\uparrow$ - - -
ChatGPT3.5 25.00 0.58 1.14 33.33 2.67 4.94 55.56 5.05 9.26 57.14 6.15 11.11 48.48 3.36 6.27
w/ CoT 45.45 2.91 5.46$\uparrow$ 33.33 2.67 4.94 40.00 4.04 7.34$\downarrow$ 47.06 6.15 10.88$\uparrow$ 43.18 3.99 7.31$\uparrow$
w/ retrieval 70.00 4.01 7.69$\uparrow$ - - - - - - 70.58 9.23 16.32$\uparrow$ - - -
GPT4 80.89 31.98 45.83 74.19 30.67 43.40 67.74 21.21 32.31 74.07 61.54 67.23 75.21 37.61 50.14
w/ CoT 86.05 21.51 34.42$\downarrow$ 73.17 40.00 51.72$\uparrow$ 80.56 29.29 42.96$\uparrow$ 71.43 76.92 74.07$\uparrow$ 75.38 41.18 53.26$\uparrow$
w/ retrieval 77.89 43.02 55.43$\uparrow$ - - - - - - 74.40 71.54 72.94$\uparrow$ - - -
Table 5: The results of CoT and retrieval techniques on the three closed-source LLMs. The $\uparrow$ and $\downarrow$ indicate whether CoT and retrieval can promote improvements in F1 score. We also provide the detection results of FaithCritic.
5 Experiments
In this section, we assess the performance of several evaluation models and specialized methods on the dataset we introduced. Thereby, we can trial the effectiveness of existing methods in detecting dialogue-level hallucination. We still conduct more fine-grained detection and explore whether the phenomenon of hallucination snowballing exists.
5.1 Baselines
We select some powerful LLMs to detect hallucination by providing specific prompts. These models include open-source LLMs: LLaMa-30B (Touvron et al., 2023), Vicuna-33B (Chiang et al., 2023), and some closed-source LLMs: Gemini1.5 PRO (Anil et al., 2023), ChatGPT3.5 (Wu et al., 2023c) and GPT4 (OpenAI, 2023). Similarly, we also experiment on specialized existing hallucination detection methods, such as FaithCritic (Dziri et al., 2022a), SelfCheckGPT (Manakul et al., 2023) and FOCUS (Zhang et al., 2023b). For a detailed description of the above baselines, please refer to Appendix A.8 (I. Baselines Selected).
5.2 Metrics
For hallucination detection, we use standard binary classification to determine whether there exists hallucination (Table 4). Meanwhile, we also conduct more fine-grained hallucination-type recognition to judge the specific subtype of hallucination and use micro-F1 score for all hallucination categories (Table 6). Appendix A.8 (II. Metrics Calculation) provides more thorough explanations.
5.3 Main Results
From the results in Table 4, we can get the following conclusions.
First, DiaHalu is a highly challenging benchmark for dialogue-level hallucination detection.
Except for GPT4, the F1 scores of all other detection methods and detecting LLMs do not exceed 50.00. Existing LLMs, such as LLaMa-30B and Vicuna-33B, are not effective in accurately discerning most samples that involve hallucination. Regarding the specialized detection methods FOCUS and SelfCheckGPT (applying prompt and NLI methods), they achieve relatively better performances. However, it proves challenging with BERTScore for SelfCheckGPT.
Second, ChatGPT3.5 shows a noticeable phenomenon of overconfidence.
Our dataset is primarily generated by ChatGPT3.5, which exhibits high confidence in its output. Despite providing a specially designed detection prompt, it still struggles to differentiate whether the dialogue content is hallucinated or not, not along the samples generated by GPT4. So, the majority of its output labels are "Non-Halu".
Third, the faithfulness hallucination is more difficult to detect for LLMs.
Apart from the specialized hallucination detection methods, the results from directly harnessing LLMs for judgment indicate that the recognition accuracy for task-oriented and chit-chat domains of dialogue are much lower than that for the knowledge-grounded and reasoning dialogue. This is because the hallucination types in the knowledge-grounded and reasoning dialogue are primarily Non-factual and Reasoning Error, which present in a more direct and apparent manner. Nevertheless, task-oriented and chit-chat domains mainly consist of three subtypes of faithfulness hallucination, which requires a LLM to possess long-term memory and the ability to recognize topics/roles transition in dialogue.
NF Ic Ir Ov RE ALL
Gemini1.5 PRO 18.97 30.49 11.36 4.76 45.41 26.72
ChatGPT3.5 1.16 4.26 0.00 0.00 9.66 3.93
GPT4 29.38 25.00 5.71 4.65 55.66 32.30
Table 6: Fine-grained hallucination-type recognition F1 scores for three LLMs. ‘NF’, ‘Ic’, ‘Ir’, ‘Ov’ and ‘RE’ stand for Non-factual, Incoherence, Irrelevance, Overreliance and Reasoning Error, respectively. ‘ALL’ represents micro-f1 of all hallucination subtypes.
Figure 5: The proportions of the three dialogue round categories. For example, the three values of R7 denote the proportions of ‘these three categories in the 7th round’ within ‘hallucinated samples that have at least seven rounds dialogues’.
5.4 Chain-of-Thought and Retrieval for Detection
Chain-of-Thought (CoT) and Retrieval are two important techniques for enhancing the ability of LLMs. In this section, we test whether these two techniques can improve the effectiveness of hallucination detection in Table 5. More details are depicted in Appendix A.11.
The experimental results indicate that both methods have facilitating effects on hallucination detection. However, Gemini1.5 PRO and ChatGPT3.5 with CoT show a decrease of around 1.00 F1 points in the chit-chat domain. We believe that these two models inherently lack the ability to recognize faithfulness hallucination, and additional CoT contents introduce noise to their judgments.
5.5 Fine-grained Hallucination-type Recognition
Table 6 shows fine-grained hallucination-type recognition results for three open-source LLMs. We can conclude that ChatGPT3.5 fails the recognition of almost all labels. To some extent, Gemini1.5 PRO and GPT4 have the ability to recognize factuality hallucination and reasoning errors, but they have lower F1 scores for the three subtypes of faithfulness hallucination. This result reveals that faithfulness hallucination remains a pressing issue for LLMs.
5.6 Hallucination Snowballing
In this section, we study the hallucination snowballing phenomenon (Zhang et al., 2023c) in our benchmark. Specifically, for each round of dialog (2-10) in all hallucinated samples, we define three categories: I hallucination that appears in previous rounds and also appears in the current round (Halu&Halu), II hallucination that appears in previous rounds but not appear in the current round (Halu&None) and III hallucination that not appear in previous rounds but appears in the current round (None&Halu). We calculate the proportions of these three categories in Figure 5.
First, I is greater than the other two categories (II and III), which means that hallucinated contents are more likely to generate new hallucinated responses. Second, I shows the most obvious increasing trend, indicating that the probability of hallucination increases with the number of dialogue rounds. These two findings validate the hallucination snowballing phenomenon.
6 Conclusion
In this paper, we propose a dialogue-level hallucination evaluation benchmark named DiaHalu. We construct the benchmark in a three-step process. The DiaHalu covers four multi-turn dialogue domains and five hallucination subtypes. Experiments through some well-known LLMs and specialized detection methods on the benchmark show that it is a challenging task, holding significant value for further research (Appendix A.12).
Limitations
This paper proposes a novel dedicated dialogue-level hallucination detection evaluation benchmark named DiaHalu. The benchmark covers four multi-turn dialogue domains and five hallucination subtypes. There is significant value for further research. However, two main limitations also exist. (1) During the second step of the benchmark construction phase, aligning the contents of speaker A with human language consumes a considerable amount of time and effort. Frequent calls to the ChatGPT3.5 or GPT4 API Keys result in a significant expense. Simultaneously, achieving consistency among all annotators led to prolonged discussion time and money cost. (2) We do not partition the dataset into training, validation, and test sets. The primary purpose of evaluation benchmarks is to assess a models’ capabilities. However, if we divide the dataset into the above three categories, this is about assigning capabilities to models. From the perspective of the two objectives, there is a clear difference. Another reason is that we need to consider the black-box detection scenario for those closed-source LLMs. However, if a division into these three types of datasets is necessarily required, it would require more data samples and larger resource consumption.
Ethics Statement
The benchmark is primarily generated by ChatGPT3.5 or GPT4. We obtain all the API Keys through a paid subscription. All the annotators are real people and they receive corresponding compensation and rewards. The entire process and outcomes are free from intellectual property and ethical legal disputes.
Acknowledgements
Due to the constrictions of the double-blind review mechanism, we will acknowledge all contributors and institutions as soon as our paper is accepted.
Appendix A Appendices
A.1 The Comparison with Other Benchmarks
In Table 1, we present the differences between DiaHalu and other hallucination detection benchmarks, highlighting the distinctive features that our DiaHalu is a naturally generated dialogue-level benchmark by LLMs, with various kinds of hallucination and explanations. All the compared benchmarks can be referred to in Section 2.1. HaluEval, WikiBio+, and PHD benchmark come from the paper Li et al. (2023b), Manakul et al. (2023) and Yang et al. (2023) respectively.
A.2 The Four Dialogue Domains
Knowledge-grounded dialogue
is designed for users to engage in knowledge-based dialogue with LLMs (Ghazvininejad et al., 2018). The knowledge includes world knowledge, factual knowledge, commonsense knowledge, and multi-hop web knowledge. It principally examines the accuracy of knowledge of the parameters in LLMs (Petroni et al., 2019).
Task-oriented dialogue
is in a form of human-computer interaction, intending to accomplish a user-specified task (Wen et al., 2017). This type of dialogue system focuses on understanding the users’ task requirements and utilizes a LLM to provide relevant information or perform specific tasks accordingly.
Chit-Chat dialogue
involves open-ended and non-goal dialogue (Sun et al., 2021). We provide two LLMs with personas and facilitate a dialogue between them. This approach allows for the evaluation of their memory capabilities, conversational coherence, and relevance to the information being discussed.
Reasoning dialogue
centralizes on the logical reasoning and understanding capabilities of LLMs. Following previous works (Chen et al., 2023b; Li et al., 2023a; Buszydlik et al., 2024; Grover et al., 2024; Zheng et al., 2023; Huang and Chang, 2023), we also treat reasoning errors as a kind of hallucination. We have the models discuss mathematical problems to achieve the answers (Kakarla et al., 2024).
A.3 The Distribution of the Sources for Dialogue Topics
The number and proportion of all dialogue topics in DiaHalu across the 10 topic sources are shown in Figure 6.
Figure 6: The distribution of the 10 sources for dialogue topics.
A.4 System Prompts for Dialogue Generation
In this section, we present the specific form of the system prompts for the four dialogue domains. We use ChatGPT3.5 (gpt-3.5-turbo-1106) [^2] https://platform.openai.com/docs/models/gpt-3-5-turbo and GPT4 (gpt-4-1106-preview) [^3] https://openai.com/gpt-4 with the temperature 0.1. The brief system prompts of the knowledge-grounded dialogue, task-oriented dialogue, chit-chat dialogue and reasoning dialogue are respectively presented in Figure 7, Figure 8, Figure 9 and Figure 10.
Previous works prove the LLMs’ ability to follow complex instructions (Manakul et al., 2023; Guan et al., 2023; Chen et al., 2023b; Mündler et al., 2023; Li et al., 2023b; Wang et al., 2023b; Liu et al., 2023; Li et al., 2024b), including some hallucination tasks. Thus, we reference such kinds of prompts, and then we formulate prompts for our benchmark.
Figure 7: The brief system prompts for knowledge-grounded dialogue.
Figure 8: The brief system prompts for task-oriented dialogue.
Figure 9: The brief system prompts for chit-chat dialogue.
Figure 10: The brief system prompts for reasoning dialogue.
A.5 Generation Details and Data Format
We utilize 8 ChatGPT3.5 and 2 GPT4 API Keys, generating approximately 4000 examples in total initially. We conduct a thorough examination of samples to filter the confused formats or contents not aligned with instructions (system prompts). To ensure a balanced number of positive and negative samples, we manually remove samples with highly similar semantics. Finally, the number of samples generated by ChatGPT3.5 and GPT4 are 748 and 355 respectively. The total number is 1103. The reason we use GPT4 to generate data is to ensure its competitiveness for a long time in the future.
Previous well-known hallucination benchmarks (Manakul et al., 2023; Guan et al., 2023; Chen et al., 2023b; Yang et al., 2023) contain 238, 400, 847, and 300 samples respectively, which is at the similar scale as ours. What’s more, our benchmark is at dialogue level, which contains about 7620 rounds (more than 27600 rounds initially) of interactions in total (6.9120 average rounds per dialogue as shown in Table 2). This indicates a larger volume of data compared to previous benchmarks at the sentence and passage levels.
It is also worth noting that, given that we assume the two subjects of the dialogue are A and B, both A and B are set to be either ChatGPT3.5 or GPT4, and it is not possible for one to be ChatGPT3.5 and the other to be GPT4.
We provide the specific format of one sample from the benchmark in Figure 11.
Figure 11: One sample from our benchmark DiaHalu
A.6 The Supplementary Details for Annotation
The annotators
all obtain at least a bachelor’s degree, get a high score in IELTS or TOEFL exams, and are proficient in using search engines such as Google and Bing. The annotators are all seasoned researchers in the field of linguistics and natural language processing. In addition, everyone exhibits strong collaborative and communicative skills. We also invite senior experts in the field of hallucination detection from academia and industry to engage in discussions and data checking.
The Annotation Process
To ensure the annotation quality, we perform three steps for annotation as described in Section 4.3 (Annotation process). First, each annotator labels around 50 samples for each domain. The annotators are required to label the presence of hallucination, hallucination subtypes and locations, along with the corresponding explanations. For cases of inconsistent annotation, we invite experts to provide suggestions in a discussion. After that, annotators specify the application scope of each hallucination label as needed. Then the annotators take a vote for resolving the label-inconsistency of the first 50 samples in each domain. Second, the entire dataset is annotated according to this standard. The annotators label all the rest samples and vote for the inconsistent samples, following data checks by the experts. Third, we conduct data statistics of the whole dataset.
In the first step above, the discussion is organized in the form of online meetings. Annotators provide the inconsistent-labeled samples to experts (first 50 in each domain), after which all annotators and experts agree on a time for an online meeting discussion. Experts provide suggestions, and annotators modify the application scopes of hallucination labels based on the suggestions, thus making it more reliable.
The Price
The annotation time for each sample ranges from 2 to 10 (average 6.2) minutes. We pay each annotator 0.5 US dollars for annotating a sample and pay each expert 0.5 US dollars for checking a sample. This exceeds the local average hourly wage. Through the aforementioned approach, the quality of the annotations and the value of the benchmark are ensured. Thus, we consider it is greatly contributory to propose such a benchmark.
Label Consistency
After the whole annotation process, we achieve a label matrix $\mathbf{L}\in\mathbb{R}^{N_{s}*N_{A}}$. $N_{s},N_{A}$ represent the number of dialogue samples and annotators respectively. The calculation for Fleiss’s Kappa is shown below:
$\begin{split}P_{e}&=\left(\frac{\sum_{i=1}^{N_{s}}\sum_{j=1}^{N_{A}}\chi_{\left\{0\right\}}(\mathbf{L}\left[i,j\right])}{N_{s}*N_{A}}\right)^{2}\\ &+\left(\frac{\sum_{i=1}^{N_{s}}\sum_{j=1}^{N_{A}}\chi_{\left\{1\right\}}(\mathbf{L}\left[i,j\right])}{N_{s}*N_{A}}\right)^{2}\end{split}$ (1)
$\begin{split}P_{o}&=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\frac{\left(\sum_{j}^{N_{A}}\chi_{\left\{0\right\}}(\mathbf{L}\left[i,j\right])\right)^{2}}{{N_{A}*\left(N_{A}-1\right)}}\\ &+\frac{\left(\sum_{j}^{N_{A}}\chi_{\left\{1\right\}}(\mathbf{L}\left[i,j\right])\right)^{2}-N_{A}}{N_{A}*\left(N_{A}-1\right)}\end{split}$ (2)
$Fleiss^{\prime}sKappa=\frac{P_{o}-P_{e}}{1-P_{e}}$ (3)
where $P_{o}$ and $P_{e}$ represent the relative observed agreement among annotators and the hypothetical probability of chance agreement respectively. $\chi_{y}(Y)$ is the Indicator Function, which means when the value of $Y$ is in set $y$, the whole function equals 1.
According to the above formulas, the calculated result for Fleiss’s Kappa of our benchmark is 0.8842, representing almost perfect agreement among all the annotators.
The Label Platform
We use Label Studio [^4] https://labelstud.io/ for labeling, which is an online open-source data labeling platform in the field of artificial intelligence. The annotation interface is depicted in Figure 12.
Figure 12: The annotation interface on Label Studio.
A.7 The Application Scope of Hallucination Labels
Annotating the hallucination and its subtypes in this benchmark is a very challenging task. One of the reasons is that some hallucination subtypes in edge cases are defiant to differentiate. Therefore, in the first stage of annotation, we provide detailed definitions for each hallucination subtype. Below are the results of the discussion between the experts and the annotators.
Non-factual
implies that it does not align with facts or introduce elements that do not exist in real life.
Incoherence
means there is a contradiction between one of the speakers and himself in the dialogue context, a contradiction between the two speakers (when both are declarative sentences) in the dialogue context. It also refers to factual and relevant nonsense, contextually inappropriate responses or other inconsistent errors.
Irrelevance
involves responses that are irrelevant to the dialogue topic or an irrelevance due to misunderstanding the grammar of a question. (Please note that we emphasize the use cases for both interrogative and declarative sentences.)
Overreliance
is that the LLM excessively trust in the correctness of the context, generating serious responses to statements that were inherently wrong or unanswerable (in a declarative sentence).
Reasoning Error
covers all errors within the reasoning dialogue.
A.8 Baselines and Metrics
I. Baselines Selected
Below is a detailed description of all the baselines we selected.
Random
A straightforward approach that randomly generates a label for each sample.
FaithCritic
(Dziri et al., 2022a) is one of the most effective dialog text hallucination classifiers before the era of large language models. Trained on a large-scale dialog corpus, it can output the confidence level for each classification label. Since the model’s input includes dialog-related knowledge, we use the retrieved contents as the knowledge during the experiment.
SelfCheckGPT
(Manakul et al., 2023) It is a widely used black-box hallucination detection framework. It rephrases the contents to be detected while ensuring the consistency of semantics by LLMs with different temperatures. Furthermore, it calculates the consistency between the original and the rephrased contents using five methods, thereby determining whether there exits hallucination. The indices $B$, $N$ and $P$ respectively denote scoring with BERTScore (Zhang et al., 2020), scoring with Natural Language Inference methods (He et al., 2023) and the direct judgment using prompts.
FOCUS
(Zhang et al., 2023b) is an improved version of SelfCheckGPT. It takes into account the attention scores between entity tokens, enabling more accurate classification of hallucination at both the sentence and paragraph levels.
LLaMa-30B && Vicuna-33B
They are two well-pretrained and widely deployed open-source LLM backbones [^5] https://huggingface.co/huggyllama/llama-30b, https://huggingface.co/lmsys/vicuna-33b-v1.3 (Touvron et al., 2023; Chiang et al., 2023). We provide a specially designed prompt to assist with detection. More details about this prompt are shown in Appendix A.9.
Gemini1.5 PRO
Gemini1.5 PRO [^6] https://gemini.google.com/ (Anil et al., 2023) is the latest version of the language model launched by Google. It inherits the powerful natural language processing capabilities of its predecessor and has made significant improvements in understanding and generating text. We employ the same prompt for binary detection as LLaMa-30B and Vicuna-33B do. We also create a manually prompt to assist with fine-grained recognition in Appendix A.10.
ChatGPT3.5 && GPT4
Both of these models are developed by OpenAI [^7] https://chat.openai.com/. ChatGPT3.5 marks the beginning of the era of large language models and GPT4 is currently the most powerful language model (Wu et al., 2023c; OpenAI, 2023). We employ the same prompt as LLaMa-30B and Vicuna-33B do for binary detection. And the same prompt as Gemeni1.5 PRO do is used for fine-grained recognition. The ChatGPT3.5 version is ChatGPT3.5 (gpt-3.5-turbo) [^8] https://platform.openai.com/docs/models/gpt-3-5-turbo and the GPT4 version is GPT4 (gpt-4-turbo)  [^9] https://openai.com/gpt-4.
II. Metrics Calculation
Despite we annotating the subtypes of hallucination in the dataset, achieving consistent labels even among humans requires further discussion. Therefore, similar to past hallucination detection efforts, we first focus on a binary classification task of determining the existence of hallucination. Consequently, we utilize binary classification evaluation metrics: Precission, Recall and F1. The positive label for this classification task is set as "Halu", for our main focus is testing the model’s ability to recognize hallucination.
As for more fine-grained hallucination-type recognition, We define a correct judgment as one where both the presence of hallucination and the specific subtype of hallucination are accurately identified. For all label types, we use micro-F1 score to quantify the performances of the three classification models.
A.9 The Prompt Designed for Detection
In Figure 13, we show the whole prompt specially designed for hallucination detection of the baselines: LLaMa-30B, Vicuna-33B, Gemini1.5 PRO, ChatGPT3.5, and GPT4. It is worth noting that due to the poor instruction-following ability and the disorderly output format of the LLaMa and Vicuna models, we conduct experiments in a 1-shot manner.
Previous works prove the LLMs’ ability to follow complex instructions (Manakul et al., 2023; Guan et al., 2023; Chen et al., 2023b; Mündler et al., 2023; Li et al., 2023b; Wang et al., 2023b; Liu et al., 2023; Li et al., 2024b), including some hallucination tasks. Thus, we reference those prompts that classify hallucination using LLMs, and then we formulate prompts for ours.
Figure 13: The whole prompt for hallucination detection of the baselines.
A.10 The Prompt Designed for Fine-grained Recognition
In Figure 14, we show the whole prompt manually created for fine-grained hallucination-type recognition of the three closed-source baselines: Gemini1.5 PRO, ChatGPT3.5, and GPT4.
Figure 14: The whole prompt for fine-grained hallucination-type recognition.
A.11 The Settings for CoT and Retrieval
Chain-of-Thought (CoT) (Feng et al., 2023a) describes the organized sequence of logical reasoning that unfolds during thinking. Retrieval (Gao et al., 2023; Ye et al., 2024) means retrieving relevant contents from the media to supplement external knowledge for LLMs. We employ CoT in all four domains to enhance the performance of ChatGPT3.5 and GPT4. The specific CoT is illustrated in Figure 15. Even so, only knowledge-grounded and reasoning domains are tested with retrieval via Google [^10] https://console.cloud.google.com/apis/library. This is because domains of task-oriented and chit-chat mainly involve scenarios related to daily life or virtual worlds, without specific domain knowledge as supplementary information.
Figure 15: The whole CoT for the four domains of dialogue.
A.12 Future Works
The necessity of dialogue-level
The sentence-level, passage-level and dialogue-level hallucination differ in the hallucination types and detection difficulties. We will explain this with the following examples to make it more clear.
Here is a sentence-level hallucination example from the dataset FactCHD (Chen et al., 2023c). ‘User: Can you tell me which mountain range is longer, the Alps or the Pyrenees? LLMs: The Pyrenees are longer than the Alps.’ In this sentence-level example, the LLM only responds to the user’s question with one sentence. We need to assess whether there are hallucination in the single sentence generated by the LLM.
Here is a passage-level hallucination example from the dataset WikiBio (Manakul et al., 2023). "Matthew Aylmer, 1st Baron Aylmer was an Irish soldier and colonial administrator. He was born in Dublin, the son of a barrister, and was educated at Trinity College, Dublin. … He was buried in Westminster Abbey." This passage-level example is directly generated by the LLM. We need to determine whether the passage with multiple sentences involves hallucination. In this example passage, hallucination occurs in the last sentence which provides unfactual information. Since there are interrelations or dependencies between the sentences in the passage, passage-level hallucination detection is more challenging than the sentence-level one.
The examples of dialogue-level hallucination are shown in Figure 2, which covers four domains and five hallucination types. The differences between it with sentence-level and passage-level hallucination are as follows: First, more types of hallucination occur frequently in dialogue. One possible reason for most benchmarks merely focusing on detecting factuality hallucination is that they are organized at a sentence or passage level. In this setting, faithfulness hallucinatios (including Incoherence, Irrelevance, and Overreliance) are less likely to occur. In contrast, since dialogue generation requires LLMs to have context coherence (Mishra et al., 2023), track the dialogue state (Heck et al., 2023), possess long-term memory capabilities (Zhong et al., 2024), and have the ability to recognize topic shifts (Lin et al., 2023), faithfulness hallucination (including Incoherence, Irrelevance, and Overreliance as described in lines 261-263) occur more frequently in dialogue. In our benchmark, the faithfulness hallucination mainly accounts for Task-oriented dialogue and Chit-Chat dialogue as shown in Figure 4. Second, it is more challenging to detect hallucination from dialogue than a single sentence or a passage. Since a dialogue contains multiple rounds of interactions that are interdependent, it can not determine whether the current round has hallucination merely based on the current content. For example, the Incoherence hallucination type occurs as the answer is not consistent with the previous context in the task-oriented dialogue in Figure 2. Thus, it needs to analyze the context dependency in the whole dialogue and judge coherence, relevance and reasoning correctness, spanning multiple rounds of interactions for dialogue-level hallucination detection, which is more challenging than sentence-level and passage-level detection.
Therefore, it is necessary to construct a dialogue-level hallucination evaluation benchmark to promote research in LLM studies.
Dialogue-level hallucination detection
is an important work in the future. We propose the first dedicated dialogue-level hallucination detection evaluation benchmark for LLMs and experimental results show that it is a very challenging task. Therefore, combining previous works (Chen et al., 2017; Deriu et al., 2021), developing methods based on this dataset to achieve a relatively high recognition accuracy is highly valuable.
Dialogue-level hallucination elimination
is an extension task of this work. Most existing hallucination elimination methods primarily focus on sentence-level or passage-level factuality hallucination (Luo et al., 2024). Hallucination elimination at the dialogue level not only requires models to have much parameter knowledge, but also a long-context memory capabilities, the abilities to recognize changes in topics/roles and logical transitions in the dialogue. These are helpful in addressing faithfulness hallucination. At the same time, improving the accuracy of knowledge in the LLMs’ parameters and the reasoning abilities are equally important.
Hallucination snowballing
is the phenomenon that LLMs tend to accumulate hallucination rather than self-correcting during the generation process (Zhang et al., 2023c). Some previous works validate this phenomenon (Azaria and Mitchell, 2023; Ang et al., 2023). In our benchamark, there is a noticeable issue of hallucination snowballing. Through the experimental results, we also display such phenomenon. Eliminating hallucination snowballing in LLMs is extremely urgent in the future.
The Unanswerability of LLMs
During the annotation process of this dataset, we introduce a hallucination category termed "overreliance", which represents answering unanswerable content (Slobodkin et al., 2023; Sulem et al., 2021). This phenomenon signifies that LLMs tend to trust the input provided by users. Sometimes, even when there are errors in user input, the LLMs still fail to recognize them. A few past researches explore the related areas and try to find a solution. However, this issue in the application of human-machine interaction and multi-agent scenarios still remains crucial.